Please show all work & explain your steps! pLease include all units as well, thank you!
Answer #1

a) given surface charge density, \sigma = -6*10-6 C/m2

since the \sigma is negative the direction of electric field will be directed towards the sheet, i.e, -ve x direction

E = \sigma/(2*\epsilono) = 6*10-6/(2*8.85*10-12) = 338983.0508 N/C

b) The direction of electric force will be directed away from the sheet,i.e, towards +ve x direction,

and, F= qE

acceleration, a=F/m =  qE/m = 4*10-6*33893.0508/(0.25) = 5.424 m/s2

c) direction of displacement will be in +ve x direction

initial velocity,u =0 m/s; displacement,s =0.5 m; a = 5.424 m/s2

using, s=ut + 0.5*at2

0.5 = 0*t + 0.5*5.424*t2

t = 0.43 seconds